Suppose you drive an old truck with a 32 gallon gas tank that gets 16 miles/gallon. If you are 75 miles from home and you have a quarter tank of gas, will you run out of gas before you get home? Show calculations.

How do you solve these types of problems?

Please help!
Answer #1

distance from home = 75 miles

mileage = 16 miles/gallon

So,

amount of gas required to reach home = distance / mileage

= 75 miles / 16 miles/gallon

= 4.7 gallon

Truck has 1/4 of 32 galllon gas = 1/4*32 = 8 gallon

Truck has 8 gallon of gas and 4.7 gallon is required to reach home.

No, it will not run out of gas before reaching home